# Flash3D: Feed-Forward Generalisable 3D Scene Reconstruction from a Single Image

### v1

@@ Abstract
In this paper, we propose Flash3D, a method for scene reconstruction and novel view synthesis from a single image which is both very generalisable and efficient. For generalisability, we start from a ‘foundation’ model for monocular depth estimation and extend it to a full 3D shape and appearance reconstructor. For efficiency, we base this extension on feed-forward Gaussian Splatting. Specifically, we predict a first layer of 3D Gaussians at the predicted depth, and then add additional layers of Gaussians that are offset in space, allowing the model to complete the reconstruction behind occlusions and truncations. Flash3D is very efficient, trainable on a single GPU in a day, and thus accessible to most researchers. It achieves state-of-the-art results when trained and tested on RealEstate10k. When transferred to unseen datasets like NYU it outperforms competitors by a large margin. More impressively, when transferred to KITTI, Flash3D achieves better PSNR than methods trained specifically on that dataset. In some instances, it even outperforms recent methods that use multiple views as input. Code, models, demo, and more results are available at https://www.robots.ox.ac.uk/~vgg/research/flash3d/.
@@ 1 Introduction
Second, we improve feed-forward per-pixel Gaussian splatting for monocular scene reconstruction. As noted, applied to single objects, a per-pixel reconstructor can use the reservoir of background pixels to model the hidden parts of the object, which is not possible when reconstructing a full scene. Our solution is to predict multiple Gaussians per pixel, where only the first Gaussian along each ray is encouraged to conform to the depth estimate and thus model the visible part of the scene. This is analogous to a layered representation [1, 3, 58, 76, 74, 36] and multi-Gaussian sampling in pixelSplat [9]. However, in our case Gaussians are deterministic, not limited to specific depth ranges, and the model is free to shift Gaussians off the ray to model occluded or truncated parts of the scene.
Overall, Flash3D is a simple and highly-performant monocular scene reconstruction pipeline. Empirically, we find that Flash3D can (a) render high-quality images of the reconstructed 3D scene, (b) operate on a wide range of scenes, both indoor and outdoor; and (c) reconstruct occluded regions, which would not be possible with depth estimation alone or with naïve extensions of it. Flash3D achieves state-of-the-art novel view synthesis accuracy on all metrics on the RealEstate10K [72]. More impressively the same, frozen model also achieves state-of-the-art accuracy when transferred to NYU [61] and KITTI [18] (in PSNR). Furthermore, in an extrapolation setting, our reconstructions can even be more accurate than those of binocular methods like pixelSplat [9] and latentSplat [83] that use two images of the scene instead of one, and are thus at a significant advantage.
In addition to the quality of the reconstructions, shown in Fig. 1, Flash3D is very efficient to evaluate and, most importantly, to train. For instance, we use 1/64th of the GPU resources of prior works like MINE [36]. By achieving state-of-the-art results while using modest computational resources for training, this opens the research area up to a wider range of researchers.
@@ 2 Related Work
Like our approach, monocular feed-forward reconstructors work by passing a single image of the scene through a neural network to output a 3D reconstruction directly. For scenes, the works of [76, 74, 75] and MINE [36] do so by predicting multi-plane images [72], and [85] uses neural radiance fields [62, 43]. Our method outperforms them in terms of speed and generalisation. Like our work, SynSin [84] uses a monocular depth predictor to reconstruct a scene; however, its reconstructions are incomplete and require a rendering network to improve the final novel views. In contrast, our approach outputs a high-quality 3D reconstruction which can be directly rendered with Gaussian Splatting [31]. For objects, a notable example is Large Reconstruction Model (LRM) [27], which obtains high-quality monocular reconstruction with a very large, and very expensive to train, model. The most related work to ours is Splatter Image [68] that uses Gaussian Splatting [31] for efficiency. Our approach also uses Gaussian Splatting as a representation, but does so for scenes rather than objects, which presents different challenges.
@@
A less challenging but still important case is few-view feed-forward reconstruction, where reconstruction requires two or more images taken from known viewpoints. Early examples used neural radiance fields (NeRF) [43] as 3D representation of objects [12, 25, 29, 38, 92, 51, 79] and scenes [12, 14, 88]. These methods implicitly learn to match points between views; the works of [10, 93] makes point matching more explicit. While many few-view reconstructors estimate the 3D shape of the object as an opacity field, an alternative is to directly predict new views [44, 55, 65, 66] of scenes with no explicit volumetric reconstruction, a concept pioneered by Light Field Networks [63]. Other works use instead multi-plane images from narrow baseline stereo pairs [72, 64] and few views [42, 30]. More related to our approach, pixelSplat [9], latentSplat [83] and MVSplat [11] reconstruct scenes from a pair of images. They utilise cross-view attention to efficiently share information and predict Gaussian mixtures to represent the scene geometry. Other very recent feed-forward approaches [69, 89, 95] combine LRM and Gaussian Splatting for reconstruction from a small number of images. We address monocular reconstruction instead, which is a much harder problem due to lack of geometric cues from triangulation.
@@
Iterative or optimisation-based methods reconstruct from one or more images by iteratively fitting a 3D model to them. Due to their iterative nature, and the need to render the 3D model to fit it to the data, they are generally much slower than feed-forward approaches. DietNeRF [28] regularises reconstruction using language models, RegNeRF [45] and RefNeRF [77] use handcrafted regularisers, and SinNeRF [87] uses monocular depth. RealFusion [40] uses an image diffusion model as a prior for monocular reconstruction based on slow score distillation sampling iterations [49]. Numerous follow-ups took a similar path [70, 86]. Convergence speed and robustness can be improved by using multi-view aware generators [82, 37, 41, 23, 73, 97]. Approaches like Viewset Diffusion [67] and RenderDiffusion [2] fuse 3D reconstruction with diffusion-based generation, which can reduce but not eliminate the cost of iterative generation. In contrast, our approach is feed-forward and therefore significantly faster, close to real-time (10fps). Some approaches generate novel views in a feed-forward manner, but iteratively and autoregressively, one view at a time. Examples include PixelSynth [52], extending SynSin, GeNVS [8], and Text2Room [26]. In contrast, we generate the final 3D reconstruction in a single feed-forward pass.
@@
Our method is based on monocular depth estimation [6, 4, 15, 33, 20, 98, 7, 21, 16, 34, 46, 50, 59, 56, 90], where metric or relative depth is predicted for every image pixel for a given image. By learning visual depth cues from large datasets, often with self-supervision, these approaches have demonstrated high accuracy and the capacity to generalise across datasets. While our method is agnostic to the depth predictor used, we use one of the state-of-the-art metric depth estimators, UniDepth [47], for our experiments.
@@ 3 Method
Let $I\in\mathbb{R}^{3\times H\times W}$ be an RGB image of a scene. Our goal is to learn a neural network $\Phi$ that takes as input $I$ and predicts a representation $\mathcal{G}=\Phi(I)$ of the 3D content of the scene, both in terms of 3D geometry and photometry. We first discuss the background and baseline model in Section 3.1, then introduce our layered multi-Gaussian predictor in Section 3.2, and finally discuss the use of monocular depth prediction as a prior in Section 3.2.
@@ 3.2 Monocular feed-forward multi-Gaussians
where $d=D(\bm{u})$ is the predicted depth at pixel $\bm{u}$ in depth map $D$ and $\delta_{1}=0$ is a constant. Note that, since the depth offset $\delta_{i}$ cannot be negative, this ensures that subsequent Gaussian layers are “behind” previous ones and encourages the network to model occluded surfaces. The mean of the $i$th Gaussian is then given by $\bm{\mu}_{i}=(u_{x}\,d_{i}/f,\,u_{y}\,d_{i}/f,\,d_{i})+\Delta_{i}.$ In practice, we find $K=2$ to be a sufficiently expressive representation.
@@
As we show empirically, it is important for the network to be able to model 3D content just outside its field-of-view. While the multiple Gaussian layers help in this regard, there is a particular need for additional Gaussians near the image border (e.g., for good new view synthesis when the camera retracts). To facilitate obtaining such Gaussians, the encoder $\Phi_{\text{enc}}$ starts by padding the input image and depth $(I,D)$ with $P>0$ pixels on each side, so that the outputs $\Phi_{k}(I,D)\in\mathbb{R}^{(C-1)\times(H+2P)\times(W+2P)}$ are larger than the inputs.
@@ 4 Experiments
We design our experiments to support four key findings, with each section dedicated to one finding. We begin with the most important result: cross-dataset generalisation—leveraging a monocular depth prediction network and training on a single dataset results in good reconstruction quality on other datasets (Section 4.2). Second, we establish that Flash3D serves as an effective representation for single-view 3D reconstruction by comparing against methods specifically designed for this task (Section 4.3). Third, we go as far as to show that the prior learned by single-view Flash3D is as strong as that learned by two-view methods (Section 4.4). Finally, we show via ablation studies how each design choice contributes to performance Flash3D (Section 4.5) and analyse outputs from Flash3D to gain insight into its inner workings.
@@ 4.1 Experiment settings
We compare with several state-of-the-art approaches that are designed specifically for single-view scene reconstruction, including LDI [76], Single-View MPI [74], SynSin [84], BTS [85] and MINE [36]. We include a comparison to our adaptation of Splatter Image [68] to show that our method is much better suited to general scenes, as well as a comparison to ray-wise unprojection $\mathcal{U}$ of the input colours to the locations predicted by the depth network. Finally, while not a fair comparison, we also compare with state-of-the-art two-view novel view synthesis methods, including [14], pixelSplat [9], MVSplat [11], and latentSplat [83].
@@
Table 1: Cross-Domain Novel View Synthesis. We evaluate NVS accuracy on datasets not used in training of our method. We outperform baselines which were trained on KITTI specifically. Here, cross-domain (CD) denotes that the method was not trained on the dataset being evaluated.
@@
Method CD PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CD PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
4.2 Cross-domain novel view synthesis
Datasets.
To the best of our knowledge, we are the first to report performance on feed-forward cross-domain monocular reconstruction. We consider two challenging comparisons. First, we evaluate Flash3D and the current state-of-the-art [36] on NYU, an indoor dataset that is similar in nature to RE10k, yet unseen in training. In Table 1, we observe that our method performs significantly better on this transfer experiment, despite the domain gap being relatively small. This suggests that prior works do not generalise as well as our method. Second, we compare our method on KITTI in Table 1, where it performs on par with the state-of-the-art that was trained on this dataset. Indeed, Flash3D outperforms the others with respect to PSNR, despite being trained only on an indoor dataset. This suggests that leveraging a pretrained depth network has allowed our network to learn an extremely strong shape and appearance prior that is even more accurate than learning on this dataset directly.
4.3 In-domain novel view synthesis
We perform an in-domain evaluation on RealEstate10k [72], following the same protocol as prior works [36]. We evaluate the quality of zero-shot reconstruction and compare performance on an in-domain dataset, RealEstate10k. We evaluate the quality of reconstruction through novel view synthesis metrics as that is the only ground-truth data available in this dataset. RealEstate10k evaluates the quality of reconstructions at different distances between the source and the target, as a smaller distance makes the task easier. In Table 2, we observe that we achieve state-of-the-art results on this mature benchmark across all distances between the source and the target. Further analysis in Fig. 3 reveals that our method’s reconstructions are sharper and more accurate than prior state-of-the-art [36], despite being trained on an order of magnitude fewer GPUs (1 vs. 64).
@@ 4.4 Comparison to few-view novel view synthesis
Unlike existing two-view methods that typically assess interpolation between two source views, Flash3D consistently performs extrapolation from a single view. The results are reported in Table 3. Here, Flash3D cannot outperform two-view approaches on the interpolation task, due to receiving less information. However, Flash3D surpasses all previous state-of-the-art two-view methods at view extrapolation. This highlights the utility of the multi-layer Gaussian representation of our approach at capturing and modelling unseen areas.
@@ 4.5 Ablation study and analysis
We ablate our method for in-domain and cross-domain settings in Table 4, focusing on the following questions. Q1: Is leveraging a monocular depth predictor useful in the task of reconstructing appearance and geometry of scenes? Q2: If yes, is it sufficient on its own, i.e., is learning shape and appearance parameters necessary for scene reconstruction with 3D Gaussians?
Importance of depth predictor. We remove the pretrained depth network that predicts depth $D$, instead estimating it jointly with all other parameters. First, in Table 4 we observe that this leads to a significant drop in performance compared to Flash3D, indicating that the depth network contains important cues that had already been learned. Moreover, the third row of Table 4 indicates that without the depth network, 2 layers of Gaussians per pixel performs worse than using just one layer. We hypothesise that the depth network plays an important role in avoiding local optima that were reported to limit the learning capabilities of primitive-based methods [9]. Qualitatively, the fourth column in Fig. 4 illustrates that removing the depth network makes it challenging to learn accurate geometries of walls (orange wall is bent) and object boundaries (bed has an incorrect shape).
Importance of extending beyond depth prediction. Here we remove the learned parts of our network. First, we use only one layer of Gaussians, and predict parameters $\mathcal{P}_{1}$ corresponding to the depth $D$ predicted by the pre-trained depth network. This also results in a drop in performance in Table 4. We then go further and remove the network that predicts $\mathcal{P}_{1}$, removing learning altogether. Novel views are simply rendered from source view colours backprojected with the depths $D$. This understandably drops the performance even further. In Fig. 4 we observe that these drops are due to the 1-layer method not representing occluded parts of the scene. The last column in Fig. 4 illustrates that the holes are significant when using only depth unprojection, and can be partially mitigated when learning shapes $\mathcal{P}_{1}$ due to the network being able to stretch the Gaussians at depth discontinuities.
@@
Fig. 5 analyses the mechanism through which the Gaussian layers form a full reconstruction of the scene. We visualise the depth of each of the layers, $D$ and $D+\delta_{2}$, multiplied by the opacity $\sigma_{1},\sigma_{2}$ of the corresponding Gaussians, illustrating how much they are used when rendering the scene. In Fig. 5, the more saturated the colour, the more opaque the Gaussian (black is fully transparent). We observe that the first layer has the most opaque Gaussians at object boundaries (wall, cabinet) and at complicated geometries (chair), indicating that these are the regions where the depth prediction network is the most useful. This is further supported by Fig. 4 where removing the depth network impaired reconstruction in exactly the same regions. Leveraging the depth network at object boundaries results in crisp, accurate geometries at small baselines (fourth column). Interestingly, the network ignores the depth prediction for windows, which are consistently incorrect. Next, we analyse where the second layer of Gaussians places its predictions. In the third column of Fig. 5,where the network observes a wall, the second layer of Gaussians is placed at larger depth. These Gaussians are observed when the camera motion (baseline) is large, and are shown to have a reasonable appearance in Fig. 5 (right column). The last column in Fig. 5 additionally reveals a limitation of our method. It is a deterministic, regressive model of structure and appearance, and thus produces blurry renderings in presence of ambiguity: when baselines are very large, in occluded regions or when camera moves backward. Blurriness could be reduced with additional losses (perceptual [96] or adversarial [22]). Alternatively, our method could be incorporated as conditioning within a framework similar to [8] or as the reconstructor in a diffusion-based feedforward 3D generation framework [67, 71].
Figure 6: Analysis of Gaussian allocation. Gaussians from the first layer (red) are allocated in visible parts, from the second layer (green) in occluded regions (top row, bottom right) and on windows (bottom left) and Gaussians from the padding region (blue) are revealed when camera reveals regions that were not present in the frustrum of the input camera.
RE10k – in-domain NYU – cross-domain KITTI – cross-domain
Flash3D 24.93 0.833 0.160 25.09 0.775 0.182 21.96 0.826 0.132
w/o depth net, w/ 2nd layer 23.62 0.782 0.186 23.73 0.732 0.210 - - -
w/o depth net, w/o 2nd layer 24.01 0.806 0.176 23.98 0.750 0.207 - - -

### v2

@@ Abstract
We propose Flash3D, a method for scene reconstruction and novel view synthesis from a single image which is both very generalisable and efficient. For generalisability, we start from a ‘foundation’ model for monocular depth estimation and extend it to a full 3D shape and appearance reconstructor. For efficiency, we base this extension on feed-forward Gaussian Splatting. Specifically, we predict a first layer of 3D Gaussians at the predicted depth, and then add additional layers of Gaussians that are offset in space, allowing the model to complete the reconstruction behind occlusions and truncations. Flash3D is very efficient, trainable on a single GPU in a day, and thus accessible to most researchers. It achieves state-of-the-art results when trained and tested on RealEstate10k. When transferred to unseen datasets like NYU it outperforms competitors by a large margin. More impressively, when transferred to KITTI, Flash3D achieves better PSNR than methods trained specifically on that dataset. In some instances, it even outperforms recent methods that use multiple views as input. Code, models, demo, and more results are available at https://www.robots.ox.ac.uk/~vgg/research/flash3d/.
@@ 1 Introduction
Second, we improve feed-forward per-pixel Gaussian splatting for monocular scene reconstruction. As noted, applied to single objects, a per-pixel reconstructor can use the reservoir of background pixels to model the hidden parts of the object, which is not possible when reconstructing a full scene. Our solution is to predict multiple Gaussians per pixel, where only the first Gaussian along each ray is encouraged to conform to the depth estimate, and thus model the visible part of the scene. This is analogous to a layered representation [1, 3, 62, 80, 78, 37] and multi-Gaussian sampling in pixelSplat [9]. However, in our case Gaussians are deterministic, not limited to specific depth ranges, and the model is free to shift Gaussians off the ray to model occluded or truncated parts of the scene.
Overall, Flash3D is a simple and highly performant monocular scene reconstruction pipeline. Empirically, we find that Flash3D can (a) render high-quality images of the reconstructed 3D scene, (b) operate on a wide range of scenes, both indoor and outdoor; and (c) reconstruct occluded regions, which would not be possible with depth estimation alone or with naïve extensions of it. Flash3D achieves state-of-the-art novel view synthesis accuracy in all metrics on RealEstate10K [76]. More impressively, the same frozen model also achieves state-of-the-art accuracy when transferred to NYU [65] and KITTI [19] (in PSNR). Furthermore, in an extrapolation setting, our reconstructions can even be more accurate than those of binocular methods like pixelSplat [9] and latentSplat [87] that use two images of the scene instead of one, and are thus at a significant advantage.
In addition to the quality of the reconstructions, shown in Fig. 1, Flash3D is very efficient to evaluate and, most importantly, to train. For instance, we use 1/64th of the GPU resources of prior works such as MINE [37]. By achieving state-of-the-art results while using modest computational resources for training, this opens the research area to a wider range of researchers.
@@ 2 Related Work
Like our approach, monocular feed-forward reconstructors work by passing a single image of the scene through a neural network to output a 3D reconstruction directly. For scenes, the works of [80, 78, 79] and MINE [37] do so by predicting multi-plane images [76], while BDS [89] uses neural radiance fields [66, 45]. Our method outperforms them in terms of speed and generalisation. Like our work, SynSin [88] uses a monocular depth predictor to reconstruct a scene; however, its reconstructions are incomplete and require a rendering network to improve the final novel views. In contrast, our approach outputs a high-quality 3D reconstruction which can be directly rendered with Gaussian Splatting [33]. For objects, a notable example is Large Reconstruction Model (LRM) [28], which obtains high-quality monocular reconstruction with a very large, and costly to train, model. The most related work is Splatter Image [72] that uses Gaussian Splatting [33] for efficiency. Our approach also uses Gaussian Splatting as a representation, but does so for scenes rather than objects, which presents different challenges.
@@
A less challenging but still important case is few-view feed-forward reconstruction, where reconstruction requires two or more images. Early examples used NeRFs [45] as 3D representation of objects [12, 26, 31, 40, 99, 54, 83] and scenes [12, 14, 93]. These methods implicitly learn to match points between views; the works of [10, 100] make point matching more explicit. While many few-view reconstructors estimate the 3D shape of the object as an opacity field, an alternative is to directly predict new views [46, 58, 69, 70] of scenes with no explicit volumetric reconstruction, a concept pioneered by LFNs [67]. Other works use instead multi-plane images from narrow baseline stereo pairs [76, 68] and few views [44, 32]. More related to our approach, pixelSplat [9], latentSplat [87] and MVSplat [11] reconstruct scenes from a pair of images. They utilise cross-view attention to share information and predict Gaussian mixtures to represent the scene geometry. Other very recent approaches [74, 94, 102] combine LRM and Gaussian Splatting for reconstruction from a small number of images. We address monocular reconstruction instead, which is a much harder problem due to lack of geometric cues from triangulation.
@@
Iterative or optimisation-based methods reconstruct from one or more images by iteratively fitting a 3D model to them. Due to their iterative nature, and the need to render the 3D model to fit it to the data, they are generally much slower than feed-forward approaches. DietNeRF [30] regularises reconstruction using language models, RegNeRF [47] and RefNeRF [81] use handcrafted regularisers, and SinNeRF [92] uses monocular depth. RealFusion [42] uses an image diffusion model as a prior for monocular reconstruction based on slow score distillation sampling iterations [52]. Numerous follow-ups took a similar path [73, 91]. Convergence speed and robustness can be improved by using multi-view aware generators [86, 39, 43, 24, 77, 104]. Approaches like Viewset Diffusion [71] and RenderDiffusion [2] fuse 3D reconstruction with diffusion-based generation, which can reduce but not eliminate the cost of iterative generation. More recently, ZeroNVS [59], ReconFusion [90] and Cat3D [17] utilize the large-scale diffusion model for single view or sparse view 3D reconstruction and NVS. However, they focus on “realistic” view generation, instead of “accurate” 3D reconstruction, and they still need the per-scene optimisation to achieve consistent 3D structure, which is relative expensive and slow. In contrast, our approach is feed-forward and therefore significantly faster, close to real-time (10fps). Some approaches generate novel views in a feed-forward manner, but iteratively and autoregressively, one view at a time. Examples include PixelSynth [55], GeNVS [8], and Text2Room [27]. In contrast, we generate the final 3D reconstruction in a single feed-forward pass.
@@
Our method is based on monocular depth estimation [6, 4, 15, 35, 21, 105, 7, 22, 16, 36, 48, 53, 63, 60, 96], where metric or relative depth is predicted for every image pixel for a given image. By learning visual depth cues from large datasets, these approaches have demonstrated high accuracy and the capacity to generalise across datasets. While our method is agnostic to the depth predictor used, we use one of the state-of-the-art metric depth estimators, UniDepth [49], for our experiments.
@@ 3 Method
Let $I\in\mathbb{R}^{3\times H\times W}$ be an RGB image of a scene. Our goal is to learn a neural network $\Phi$ that takes as input $I$ and predicts a representation $\mathcal{G}=\Phi(I)$ of the 3D content of the scene, both in terms of 3D geometry and photometry. We first discuss the background and baseline model in Sec. 3.1, introduce our layered multi-Gaussian predictor and discuss the use of monocular depth prediction as a prior in Sec. 3.2.
@@ 3.2 Monocular feed-forward multi-Gaussians
where $d=D(\bm{u})$ is the predicted depth at pixel $\bm{u}$ in depth map $D$ and $\delta_{1}=0$ is a constant. Note that since the depth offset $\delta_{i}$ cannot be negative, this ensures that subsequent Gaussian layers are “behind” previous ones and encourages the network to model occluded surfaces. The mean of the $i$th Gaussian is then given by $\bm{\mu}_{i}=(u_{x}\,d_{i}/f,\,u_{y}\,d_{i}/f,\,d_{i})+\Delta_{i}.$ In practice, we find $K=2$ to be a sufficiently expressive representation.
@@
As we show empirically, it is important for the network to be able to model 3D content just outside its field-of-view. Although multiple Gaussian layers help in this regard, there is a particular need for additional Gaussians near the image border (e.g., for good new view synthesis when the camera retracts). To facilitate obtaining such Gaussians, the encoder $\Phi_{\text{enc}}$ starts with padding the input image and depth $(I,D)$ with $P>0$ pixels on each side, so that the outputs $\Phi_{k}(I,D)\in\mathbb{R}^{(C-1)\times(H+2P)\times(W+2P)}$ are larger than the inputs.
@@ 4 Experiments
We design our experiments to support four key findings, with each section dedicated to one finding. We begin with the most important result: cross-dataset generalisation—leveraging a monocular depth prediction network and training on a single dataset results in good reconstruction quality on other datasets (Sec. 4.2). Second, we establish that Flash3D serves as an effective representation for single-view 3D reconstruction by comparing against methods specifically designed for this task (Sec. 4.3). Third, we go as far as to show that the prior learned by single-view Flash3D is as strong as that learned by two-view methods (Sec. 4.4). Finally, we show via ablation studies how each design choice contributes to performance Flash3D (Sec. 4.5) and analyse outputs from Flash3D to gain insight into its inner workings.
@@ 4.1 Experiment settings
We compare with several state-of-the-art single-view 3D scene reconstruction models, including LDI [80], Single-View MPI [78], SynSin [88], BTS [89] and MINE [37]. We include a comparison to our adaptation of Splatter Image [72] to show that our method is much better suited to general scenes, as well as a comparison to ray-wise unprojection $\mathcal{U}$ of the input colours to the locations predicted by the depth network. Finally, while not a fair comparison, we also compare with state-of-the-art two-view novel view synthesis methods, including [14], pixelSplat [9], MVSplat [11], and latentSplat [87].
@@
Table 1: Cross-Domain Novel View Synthesis. We evaluate Novel View Synthesis accuracy on datasets not used in training of our method. We outperform baselines which were trained on KITTI specifically. Here, cross-domain (CD) denotes that the method was not trained on the dataset being evaluated.
@@
[…]
@@
4.2 Cross-domain novel view synthesis
Datasets.
To the best of our knowledge, we are the first to report performance on feed-forward cross-domain monocular reconstruction. We consider two challenging comparisons. First, we evaluate Flash3D and the current state-of-the-art [37] on NYU, an indoor dataset that is similar in nature to RE10k, yet unseen in training. In Tab. 1, we observe that our method performs significantly better on this transfer experiment, despite the domain gap being relatively small. This suggests that prior works do not generalise as well as our method. Second, we compare our method on KITTI in Tab. 1, where it performs on par with the state-of-the-art that was trained on this dataset. Indeed, Flash3D outperforms the others with respect to PSNR, despite being trained only on an indoor dataset. This suggests that leveraging a pretrained depth network has allowed our network to learn an extremely strong shape and appearance prior that is even more accurate than learning on this dataset directly.
RE10k – in-domain NYU – cross-domain KITTI – cross-domain
Flash3D 24.93 0.833 0.160 25.09 0.775 0.182 21.96 0.826 0.132
w/o depth net, w/ 2nd layer 23.62 0.782 0.186 23.73 0.732 0.210 - - -
w/o depth net, w/o 2nd layer 24.01 0.806 0.176 23.98 0.750 0.207 - - -
@@ 4.2 Cross-domain novel view synthesis
4.3 In-domain novel view synthesis
We perform an in-domain evaluation on RealEstate10k [76], following the same protocol as prior works [37]. We evaluate the quality of zero-shot reconstruction and compare performance on an in-domain dataset, RealEstate10k. We evaluate the quality of reconstruction through novel view synthesis metrics as that is the only ground-truth data available in this dataset. RealEstate10k evaluates the quality of reconstructions at different distances between the source and the target, as a smaller distance makes the task easier. In Tab. 2, we observe that we achieve state-of-the-art results on this mature benchmark across all distances between the source and the target. Further analysis in Fig. 3 reveals that our method’s reconstructions are sharper and more accurate than prior state-of-the-art [37], despite being trained on an order of magnitude fewer GPUs (1 vs. 64).
@@ 4.4 Comparison to few-view novel view synthesis
Unlike existing two-view methods that typically assess interpolation between two source views, Flash3D consistently performs extrapolation from a single view. The results are reported in Tab. 4. Here, Flash3D cannot outperform two-view approaches on the interpolation task, due to receiving less information. However, Flash3D surpasses all previous state-of-the-art two-view methods at view extrapolation. This highlights the utility of the multi-layer Gaussian representation of our approach at capturing and modelling unseen areas.
@@ 4.5 Ablation study and analysis
We ablate our method for in-domain and cross-domain settings in Tab. 3, focusing on the following questions. Q1: Is leveraging a monocular depth predictor useful in the task of reconstructing appearance and geometry of scenes? Q2: If yes, is it sufficient on its own, i.e., is learning shape and appearance parameters necessary for scene reconstruction with 3D Gaussians?
Importance of depth predictor. We remove the pretrained depth network that predicts depth $D$, instead estimating it jointly with all other parameters. First, in Tab. 3 we observe that this leads to a significant drop in performance compared to Flash3D, indicating that the depth network contains important cues that had already been learned. Moreover, the third row of Tab. 3 indicates that without the depth network, 2 layers of Gaussians per pixel performs worse than using just one layer. We hypothesise that the depth network plays an important role in avoiding local optima that were reported to limit the learning capabilities of primitive-based methods [9]. Qualitatively, the fourth column in Fig. 4 illustrates that removing the depth network makes it challenging to learn accurate geometries of walls (orange wall is bent) and object boundaries (bed has an incorrect shape).
Importance of extending beyond depth prediction. Here we remove the learned parts of our network. First, we use only one layer of Gaussians and predict parameters $\mathcal{P}_{1}$ corresponding to the depth $D$ predicted by the pre-trained depth network. This also results in a drop in performance in Tab. 3. We then go further and remove the network that predicts $\mathcal{P}_{1}$, removing learning altogether. Novel views are simply rendered from source view colours backprojected with the depths $D$. This understandably drops the performance even further. In Fig. 4 we observe that these drops are due to the 1-layer method not representing occluded parts of the scene. The last column in Fig. 4 illustrates that the holes are significant when using only depth unprojection, and can be partially mitigated when learning shapes $\mathcal{P}_{1}$ due to the network being able to stretch the Gaussians at depth discontinuities.
@@
Fig. 5 analyses the contribution of each Gaussian layer to a full reconstruction of the scene. We visualise the depth of each of the layers, $D$ and $D+\delta_{2}$, multiplied by the opacity $\sigma_{1},\sigma_{2}$ of the corresponding Gaussians, illustrating how much they are used when rendering the scene. In Fig. 5, the more saturated the colour, the more opaque the Gaussian (black is fully transparent). We observe that the first layer has the most opaque Gaussians at object boundaries (wall, cabinet) and at complicated geometries (chair), indicating that these are the regions where the depth prediction network is the most useful. This is further supported by Fig. 4 where removing the depth network impaired reconstruction in exactly the same regions. Leveraging the depth network at object boundaries results in crisp, accurate geometries at small baselines (fourth column). Interestingly, the network ignores the depth prediction for windows, which are consistently incorrect. Next, we analyse where the second layer of Gaussians places its predictions. In the third column of Fig. 5,where the network observes a wall, the second layer of Gaussians is placed at larger depth. These Gaussians are observed when the camera motion (baseline) is large, and are shown to have a reasonable appearance in Fig. 5 (right column). The last column in Fig. 5 additionally reveals a limitation of our method. It is a deterministic, regressive model of structure and appearance, and thus produces blurry renderings in presence of ambiguity: when baselines are very large, in occluded regions or when camera moves backward. Blurriness could be reduced with additional losses (perceptual [103] or adversarial [23]). Alternatively, our method could be incorporated as conditioning within a framework similar to [8] or as the reconstructor in a diffusion-based feed-forward 3D generation framework [71, 75].